OpenStudy (anonymous):

I'll become a fan and give a medal to the first person to answer: A regular pentagonal pyramid with base edge 1.5 and slant height 9. What is the Lateral area?

Directrix (directrix):

The lateral faces of a regular pyramid are congruent isosceles triangles. The altitude of any of these triangles is the slant height of the regular pyramid. The lateral area of a regular pyramid is given by the formula: LA = (1/2)* p * L where p is the perimeter of the base of the pyramid and L is the slant height of the pyramid. LA = (1/2) * (5*1.5) * 9 LA = ? @Mrfootballman97